OpenStudy (anonymous):

what are the roots of the equation z^2+11z-42

OpenStudy (anonymous):

-14 and 3 are the roots

OpenStudy (anonymous):

but when added its 11 not -11

OpenStudy (anonymous):

i figured it out its -3, 14

OpenStudy (mertsj):

No it isn't.

OpenStudy (anonymous):

what ever it may be... you got it or not....

OpenStudy (anonymous):

yes it is -3+14=11 -3*14=-42

OpenStudy (mertsj):

In the first place z^2+11z-42 is not an equation

OpenStudy (mertsj):

So what do you mean when you say what are the roots of the equation? What equation?

OpenStudy (anonymous):

yes then (z+14)(z-3)=0 then z=-14 & z=3

OpenStudy (anonymous):

yes it is. Its in my text book

OpenStudy (anonymous):

nope

OpenStudy (mertsj):

Does it have an equal sign?